Please help me get the equation for this problem:
If the length of each side of a square is decreased by 4in., the perimeter of the new square is 10 in. more than half the perimeter of the original square. What are the dimension of the original square?

Answer by nerdybill(7384)

If the length of each side of a square is decreased by 4in., the perimeter of the new square is 10 in. more than half the perimeter of the original square. What are the dimension of the original square?
Let x = length of one side of original square
then
x-4 = length of side decreased by 4
.
perimeter of original square = 4x
perimeter of new square = 4(x-4) = 4x-16
.
From:
"perimeter of the new square is 10 in. more than half the perimeter of the original square" we get our equation:
4x-16 = (1/2)(4x) + 10
4x-16 = 2x + 10
2x-16 = 10
2x = 26
x = 13
.
What are the dimension of the original square?
13 inches by 13 inches